OpenStudy (anonymous):

Given a = -3, b = 4 and c = -5, evaluate -|a - b| + c.

OpenStudy (anonymous):

-|(-3) - (4)| + (-5) -|(-3) - (4)| = -7 -> 7 - abs(7) = -7 -7 + _5 = -7-5 = -12